Find an equation in slope-intercept form for the line with x-intercept -2/3 and perpendicular to 2x-y=4.
Answer by stanbon(75887)

Find an equation in slope-intercept form for the line with x-intercept -2/3 and perpendicular to 2x-y=4.
-------
Find the slope of the given line:
y = 2x-4
slope = 2
---
Any line perpendicular to it must have slope = -1/2
----------------------------------
Form of the equation: y = mx + b
Solve for "b":
0 = (-1/2)(-2/3) + b
0 = (1/3) + b
b = -1/3
-----
Equation:
y = (-1/2)x -(1/3)
=========================
Cheers,
Stan H.
================